# Aperiodic points for outer billiards

### v3

@@ 7. Summary and perspectives
Let $\Pi$ be a regular $N$-gon with $N\neq 3$, $4$, $5$, $6$. A theorem of Rukhovich (Theorem 3.1) allows to assume that $\Pi$ has an even number of vertices, equivalently, that $\Pi$ is centrally symmetric. By the results of Section 3.1, the polygonal domain $X$ defined there is invariant under $f_{\Pi}$; the restriction of $f_{\Pi}$ to $X$ is denoted by $f_{X}$. The central symmetry $M:z\mapsto-z$ of $\Pi$ is also a symmetry of $X$, therefore, $g_{X}:=M\circ f_{X}$ is a polygon exchange transformation of $X$. For $g_{X}$, it is not hard to compute the dynamical Hadwiger valuation (Theorem 3.6). Since this valuation does not vanish (Corollary 3.7), it follows (Theorem 2.29) that $g_{X}$ is not periodic. Further, Theorem 6.1 implies the existence of aperiodic points for $g_{X}$. Also, it is easy to see that $f_{X}$ has aperiodic points if and only if $g_{X}$ does — it suffices to note that $M$ commutes with $f_{X}$. In this way, we obtain the result of Theorem 1.1.
@@ 7.2. Comparison of the methods
In Sections 3 and 4, we gave two methods for establishing non-periodicity of certain scissors automorphisms. Method 1 uses dynamic Hadwiger invariants. For odd $N$, one can directly compute the Hadwiger dynamical invariant of $f_{X}^{2}$ (which acts by translations) — the only problem is that the computation gets technically more involved. An easier computation yields the value of the invariant for $g_{X}$, and, in fact, general properties of dynamical valuations imply that $\mathrm{Inv}_{L}(f_{X}^{2})=2\mathrm{Inv}_{L}(g_{X})$, for any choice of a co-oriented line $L$. Computing dynamic Hadwiger invariants for more general polygon exchange transformations may not be as easy. Note also that one needs to reduce a given piecewise isometry to a $\mathsf{T}$-scissors automorphisms in order to use Hadwiger invariants.
Method 2 was used to prove a special case of Theorem 3.8, namely, the case of a twice odd $N$. However, we believe that this method is also applicable to arbitrary regular polygons $\Pi$, possibly at the expense of more significant technical details. Similarly to method 1, the second method uses dynamic valuations, but this time in dimension one. The first step is therefore to reduce some properties of 2D dynamics to relevant properties of 1D dynamics — the notion of a boundary quasi-commutator and Theorem 4.3 serve this purpose. If a boundary quasi-commutator of a given map has nonzero SAF invariant, then the map cannot be periodic. Computations are simpler in the case when $N$ is twice odd and, for this reason, we worked out only this case (see Theorem 4.13); there are no principal obstructions to extending the computations to all $N$. From a general perspective, method 2 looks more robust: firstly, it applies to piecewise rotations, not only to piecewise translations, and secondly, it operates with SAF invariants of interval exchange transformations, which are better understood that multi-dimensional analogs.
@@ 7.3. Possible extensions
In situation of Theorem 1.1, dynamical valuations can be computed explicitly. One possible direction for further research is to extend Theorem 1.1 to all non-lattice quasi-rational polygons (a conjecture of Gutkin and Simanyi). An additional difficulty of this setup is that it is too general for relying on closed form formulas. General methods will be necessary for establishing the nonvanishing of dynamic valuations.
Related is a problem of better understanding what exactly nonzero dynamic valuations can say about dynamics, besides the lack of periodicity. This problem is not fully resolved even in dimension one, that is, for interval exchange transformations. It would be natural to first approach this problem either for special classes of interval exchange maps or for outer billiards on regular polygons.
@@ 7.4. Further questions
  • Find interesting classes of multi-dimensional polytope exchange transformations, for which dynamic valuations are explicitly computable.
  • Relate polytope exchange transformations in dimension $d$ with flows on Euclidean $(d+1)$-manifolds; describe the role of dynamic valuations in terms of these flows.
  • Develop general methods for establishing (non)existance of dynamic self-similarity in the context of piecewise rotations.

### v4

@@ 7. Summary and perspectives
Let $\Pi$ be a regular $N$-gon with $N\neq 3$, $4$, $5$, $6$. A theorem of Rukhovich (Theorem 3.1) allows to assume that $\Pi$ has an even number of vertices, equivalently, that $\Pi$ is centrally symmetric. By the results of Section 3.1, the polygonal domain $X$ defined there is invariant under $f_{\Pi}$; the restriction of $f_{\Pi}$ to $X$ is denoted by $f_{X}$. The central symmetry $M:z\mapsto-z$ of $\Pi$ is also a symmetry of $X$, therefore, $g_{X}:=M\circ f_{X}$ is a polygon exchange transformation of $X$. For $g_{X}$, it is not hard to compute the dynamic Hadwiger invariant (Theorem 3.6). Since this valuation does not vanish (Corollary 3.7), it follows (Theorem 2.29) that $g_{X}$ is not periodic. Further, Theorem 6.1 implies the existence of aperiodic points for $g_{X}$. Also, it is easy to see that $f_{X}$ has aperiodic points if and only if $g_{X}$ does — it suffices to note that $M$ commutes with $f_{X}$. In this way, we obtain the result of Theorem 1.1.
@@ 7.2. Comparison of the methods
In Sections 3 and 4, we gave two methods for establishing non-periodicity of certain scissors automorphisms. Method 1 uses dynamic Hadwiger invariants. For odd $N$, one can directly compute the dynamic Hadwiger invariant of $f_{X}^{2}$ (which acts by translations) — the only problem is that the computation gets technically more involved. An easier computation yields the value of the invariant for $g_{X}$, and, in fact, general properties of dynamic invariants imply that $\mathrm{Inv}_{L}(f_{X}^{2})=2\mathrm{Inv}_{L}(g_{X})$, for any choice of a co-oriented line $L$. Computing dynamic Hadwiger invariants for more general polygon exchange transformations may not be as easy. Note also that one needs to reduce a given piecewise isometry to a $\mathsf{T}$-scissors automorphisms in order to use Hadwiger invariants.
Method 2 was used to prove a special case of Theorem 3.8, namely, the case of a twice odd $N$. However, we believe that this method is also applicable to arbitrary regular polygons $\Pi$, possibly at the expense of more significant technical details. Similarly to method 1, the second method uses dynamic invariants, but this time in dimension one. The first step is therefore to reduce some properties of 2D dynamics to relevant properties of 1D dynamics — the notion of a boundary quasi-commutator and Theorem 4.3 serve this purpose. If a boundary quasi-commutator of a given map has nonzero SAF invariant, then the map cannot be periodic. Computations are simpler in the case when $N$ is twice odd and, for this reason, we worked out only this case (see Theorem 4.13); there are no principal obstructions to extending the computations to all $N$. From a general perspective, method 2 looks more robust: firstly, it applies to piecewise rotations, not only to piecewise translations, and secondly, it operates with SAF invariants of interval exchange transformations, which are better understood than their multi-dimensional analogs.
@@ 7.3. Possible extensions
In situation of Theorem 1.1, dynamic invariants can be computed explicitly. One possible direction for further research is extending Theorem 1.1 to all non-lattice quasi-rational polygons (a conjecture of Gutkin and Simanyi). An additional difficulty of this setup is that it is too general for relying on closed form formulas. General methods will be necessary for establishing the nonvanishing of dynamic invariants.
Related is a problem of better understanding what exactly nonzero dynamic invariants can say about dynamics, besides the lack of periodicity. This problem is not fully resolved even in dimension one, that is, for interval exchange transformations. It would be natural to first approach this problem either for special classes of interval exchange maps or for outer billiards on regular polygons.
@@ 7.4. Further questions
  • Find interesting classes of multi-dimensional polytope exchange transformations, for which dynamic invariants are explicitly computable.
  • Relate polytope exchange transformations in dimension $d$ with flows on Euclidean $(d+1)$-manifolds; describe the role of dynamic invariants in terms of these flows.
  • Develop general methods for establishing (non)existence of dynamical self-similarity in the context of scissors automorphisms.